OpenStudy (anonymous):

what is the first derivate of x(t)= Et^2 - Ft + G

OpenStudy (anonymous):

E and F and G would be constants just like numbers.

OpenStudy (rizags):

split it up into derivative of each term

OpenStudy (rizags):

differentiate each term individually

OpenStudy (anonymous):

I had Et + F + G but i was told it is wrong

OpenStudy (rizags):

do you know what \[\frac{d}{dt}G\] is, where G is a constant?

OpenStudy (anonymous):

it becomes zero right ?

OpenStudy (rizags):

look, all you need to do is this: \[\frac{d}{dt}Et^2-\frac{d}{dt}Ft+\frac{d}{dt}G\]

OpenStudy (rizags):

and yes

OpenStudy (anonymous):

How did you get +F from a -Ft? It would be a -F instead. Derivative of Et^2 is not Et Use power rule Derivative of \[ax^n\] where a is a constant is \[anx ^{n-1}\]

OpenStudy (anonymous):

so the answer should be E 2t - F

OpenStudy (rizags):

\[\frac{dx}{dt} x(t)=Et^2-Ft+G=2Et-F\]

OpenStudy (anonymous):

2ET would be better than E2t. Always good to have the number in front.